A closed-end investment company has a net asset value of $12.00. A year ago the shares sold for a 16 percent discount but that discount has narrowed (that is, declined) to 8 percent. If the company distributed $1.25 a share, what was the return on an investment in the shares before considering commissions on the purchase? Round your answer to two decimal places. %
Answer #1

current price = NAV * (1- discount)

= 12 * (1-0.16)

= 10.08

selling price = 12 * (1-0.08)

= 11.04

return = P1+D-P0/Po

= 11.04+1.25-10.08/10.08

= 21.93%